### v1

@@ 2.1 Clustering Based on Visits
We hypothesize that users may differ in terms of their exercise habits. These differences are often attributed to age and gender; however, external factors like exercise environment and daily routines also determine when people find opportunities for these activities. We used Non-negative Matrix Factorization (NMF) to identify distinct behavioral groups when studying daily and weekly gym visits. The idea here was that different groups may have varying tendencies and independent habit formations, and we could incorporate the group-specific differences. The detailed explanations regarding the methodology are presented in Section 4.2. The clustering led us to the five apparent clusters in all clubs, shown in Figure 1(a). Firstly, there is the morning cluster, who visit the gyms predominantly around 8-9 AM. Then, there is the noon cluster, who go to the gym around 12-13. After that, there is the afternoon cluster, who tend to visit the most approximately between 15-17. Next is the evening cluster who go to the establishments around 18-19. Finally, there is the night cluster, who like to visit the gym when it is mostly dark outsize, around 20-21. We notice morning and night clusters exhibit different patterns for weekend days, because the gyms operate between the hours 6-23 on the weekdays and 8-20 on the weekends.
We classified and named these behavioral groups according to the hours that they most visited. The density of the visits presented in the heatmap to reflect the most preferred times for each cluster. The first 6 weeks of visitation patterns for each user is used for clustering, but we show the assigned clusters are consistent when compared to 17th week.
@@
Figure 1: Behavioral clusters and preferences by demographics. Non-negative matrix factorization points to 5 distinct exercise groups. These groups differ mainly by their preference to exercise during the day (a). Some of these clusters are more preferred by some gender and age groups (b). Since the cluster assignments were determined by user activities in the first 6 weeks, we measured the confidence of the NMF model for different models (c,d). Finally, we check the consistency of assignment when data for the first 6 and 17 weeks are considered (e).
After we established these clusters, we wanted to find out how much each cluster deviated from our overall customer in multiple facets. For that, we looked at the gender and age groups and their occurrence rate in particular clusters. Figure 1(b) shows the deviations from the overall behavior for that group. The deviation measures how much more or less likely to observe particular type of users within that cluster. We calculate deviation as the difference in probabilities that we calculate for different groups. As an example, 0.01 deviation for females in the first subfigure in Fig1(b) implies that in the Morning cluster, it is 1% more likely to find a woman than overall. Conversely, we are 53% less likely to pick a person between the ages 14-20 in the Morning cluster, depicted in the same subfigure. Detailed methodology for the calculations are available in Section 4.2.
@@
Younger people (ages 14-27) tend to do less visits in the morning and noon, as indicated by their negative deviation in the two upper subfigures in Fig1(b). Older people (ages 49+); however, make more visits in the earlier parts of the day, as their morning deviations are in the positive, and the rest are in the negative. This can be attributed to visiting just before work to allow themselves the proper free time after work. Their deviations also stand out, which we associate with less number of people in those groups. Women are more prominent in the noon and afternoon clusters, while men, as they make up the majority of people, do not deviate too much, but visit more often in the afternoon and at night. Most starkly, the people in the ages of 14-20 make most of their visits in the afternoon. This may be due to their schedules allowing them to visit the gym after school/university.
@@
Figures 1(c) and (d) show the cumulative distribution function for the probability of belonging to a cluster for each different cluster, signified by their unique color. Looking at these figures, it is clear that the NMF decomposition produces more confident results for the morning and night clusters, while the probabilities produced for the people in the rest of the clusters are less confident. This is also evident by the heatmaps produced in Fig1(a), as there are a lot of overlaps for the rest of the clusters, while the morning and night clusters are more separated. We theorize that people that work out in the morning and at night have more fixed schedules, preventing them from visiting at different hours. The rest of the clusters have more flexible schedules, allowing them to change their gym hours more frequently, thus making them harder to assign to a single cluster.
Figure 1(e) presents the cluster transition matrix for the important weeks of 6 and 17. On both axes, we have the 5 clusters, and the annotations represent some percent of the population. Although the previous two figures showed that morning and night clusters were more confidently assigned than the rest, this figure proves that clusters assignments are robust for an extended period of time, as the diagonal values account for more than 84% of the population, which indicates that people do not transition between different behavioral clusters. Additionally, most of the transitions are from clusters with closer times, which is to be expected.
2.2 Proposition on a Measurement of Habit Formation
Understanding the underlying mechanisms of habit formation is key to developing targeted interventions that motivate individuals to attend the gym consistently. We developed the survival metric to measure and analyze the persistence of gym habits. Our approach to this analysis was inspired by the work of Harris and Kessler (2019), who explored habit formation and activity persistence in the context of gym equipment usage. Their study demonstrated that frequent early activity leads to more persistent exercise behavior, suggesting that interventions aiming for behavioral change need to involve higher frequency or longer duration activities to be effective [25]. This metric was defined as the number of consecutive weeks a member went to the gym at least once. We added a tolerance of one week of absence called gap week, recognizing that occasional interruptions (e.g., travel or illness) should not disrupt habit formation. The rationale behind creating the survival metric was to ensure that habit attendance could be measured. This metric could then be used to identify factors that influence the habit formation process.
@@ 2.2 Proposition on a Measurement of Habit Formation
Our analysis of the survival metric revealed clear patterns in gym attendance behavior and allowed us to identify essential milestones that indicate the progression of habit formation. As shown in Figure 2(a), the cumulative distribution of members’ survival series shows a rapid increase in gym attendance in the first weeks. This rapid increase may reflect high initial motivation. However, we observed that approximately 50% of members did not maintain their attendance streak at the 6-week point. This 6-week milestone emerged as a critical point in the habit formation process where members either solidify or begin to abandon their gym habit. When we extended our analysis to users who survived even longer, we observed a steady decline in the number of members maintaining their attendance streak. At the 17-week point, only 20% of the member population maintained their streak, indicating an increasing difficulty in maintaining long-term attendance. Members with a survival streak between these two weeks are habit-holders.
@@
To further assess and track habit formation, we established specific attendance milestones corresponding to critical survival durations. These milestones mark a member’s progress in their gym attendance journey. Each milestone is defined by a critical visit count that members must achieve within the specified period. For example, to reach the $\text{6}^{\text{th}}$ week milestone, a member must attend the gym at least 9 times within the first six weeks of their membership. Figure 3(d) shows an almost linear trend of the critical visits to reach milestones. We fitted a linear model $y=2.01x-5.35$, where $m\approx 2.01$, indicating that members need to visit the gym roughly twice per week to meet their milestones. This suggests that attending the gym at least twice a week is important for maintaining consistent habits. The decision to create these milestones was due to the need for a structured approach to evaluate habit formation over time. A user can still attend gyms for an extended period of time, but we use milestones to separate users with the same length of membership into two categories. By focusing on these milestones, we aimed to simplify the tracking of member progress and to identify points at which interventions might be most effective. The details of how these milestones were achieved are discussed in Section 4.4.
@@ 2.3 Impact of Interventions on Habit Formation
Deriving meaningful insights in fields such as public health, economics, and social sciences often relies on identifying the underlying causal relationships between various interventions and outcomes. To achieve this, we implemented a causal inference model designed to estimate the impact of specific gym-related interventions on the consistency of exercise habits among gym members. We use propensity score matching (PSM) framework, which allows us to isolate the effect of particular interventions, such as participation in group lessons, personal trainer sessions, visiting different clubs, and invitation credit usage, while controlling for confounding variables that might otherwise introduce bias into the results. In our model, we first identified the treatment and control groups for each intervention. The treatment group consisted of gym members who participated in the specific intervention, while the control group comprised members who did not engage in that intervention but were otherwise similar based on matched propensity scores. Propensity scores were calculated using a variety of covariates, including demographic factors such as age and gender, self-reported experience level, the type of cluster they were in, body mass index, and membership type. These covariates were carefully selected to account for factors that could simultaneously affect both the likelihood of receiving the treatment and the outcome, which in this case is the formation of consistent exercise habits. By matching gym members in the treatment and control groups based on their propensity scores, we aimed to reduce confounding and better approximate the conditions of a randomized controlled trial.
To assess the impact of interventions on habit formation, we focused on the first six weeks of gym attendance, a period identified as critical for habit formation in Section 2.2. There are several variables within these six weeks that we use to analyze their influence on sustained exercise habits. This approach allowed us to observe how early interventions and behaviors contribute to the long-term success of habit formation. Critical visit counts for each week were accepted as the reference point for determining habit formation. By integrating critical visit measure into our model, we were able to provide a more precise estimation of how early interventions, as measured by the critical visits, affect the likelihood of members developing consistent exercise habits.
Once the groups were matched, we estimated the treatment effects by analyzing the differences in outcomes (habit formation) between the treated and control groups. The coefficients derived from this analysis represent the estimated effect size of the interventions. Positive coefficients suggest that the intervention has a beneficial impact on the development of consistent exercise habits, while negative coefficients indicate a detrimental effect. Coefficients close to zero suggest a negligible impact. This method provided us with a robust framework to evaluate the significance of early interventions in shaping long-term exercise habits.
In our study, we employed DoWhy [26], an end-to-end library for causal inference, to implement this causal model. DoWhy facilitated the identification, estimation, and validation of causal effects, allowing us to rigorously assess the impact of different gym-related interventions on habit formation. The overall effects of various gym-related interventions on habit formation, shown in Fig4(a).
@@
Personal trainer sessions exhibit a strong positive impact on maintaining regular exercise habits. The coefficients for personal training are consistently higher than those for other interventions, indicates the significant role of personalized guidance and accountability in habit formation. Personal trainers offer customized workout plans and motivational support, which appear to be significant in helping members establish and maintain their exercise routines.
@@
The results of the cluster analysis also indicates that specific interventions have diverse impacts on distinct subgroups of gym members. The impact of these interventions across different clusters, specifically highlighting the effects observed at the $\text{6}^{\text{th}}$ and $\text{17}^{\text{th}}$ weeks shown in Fig4(b).
@@
Noon cluster members exhibit a significantly higher positive response to invitation credit usage and group lessons compared to other clusters. This group tends to be younger adults, with a higher-than-average proportion of females. Since the noon cluster hours coincide with working hours in Türkiye, it is possible that noon cluster members have more flexible schedules. The flexible schedules of the noon cluster may allow them to view the gym as a social hub, where they can combine exercise with socializing, which is particularly appealing to younger members. Additionally, these members show a sustained positive effect from taking personal trainer sessions, particularly by week 17. This indicates that personalized guidance and accountability provided by personal trainers continue to support consistent exercise habits over time, making it an important intervention for this group.
@@
Evening cluster members have a higher-than-average proportion of females, which influences their response to interventions. Unlike other clusters, this group shows a lower positive response to group lessons. Although the female presence is significantly higher, similar to the noon cluster, the cluster behavior differs significantly. The overall effect of group lessons on the evening cluster is negligible, with the impact being lower than the entire population (0.18), suggesting that group workouts may not be as appealing to them. However, this cluster shows the highest and most sustained positive effect from personal trainer sessions by the $\text{17}^{\text{th}}$ week, making it the cluster that best maintains its benefits from personalized training. This indicates that for evening exercisers, particularly in a female-represented cluster, personalized attention and tailored workouts play a crucial role in fostering long-term exercise commitment, especially during the evening hours after a day’s work.
Night clusters benefits from a mix of interventions, including group lessons and personal training sessions. The late-night hours are typically suitable for those with unconventional schedules, such as students or shift workers. The fact that this group is predominantly younger, with older individuals being strongly underrepresented also supports this. Working with a personal trainer has the most significant effect on habit formation within this cluster during the first 6 weeks. However, this effect diminishes considerably by the $\text{17}^{\text{th}}$ week. This decline may suggest that younger members, who form the majority of this cluster, initially benefit from the structure and guidance provided by a personal trainer but may choose to transition to more independent workouts as they gain confidence and familiarity with their routines. The initial high impact of personal training likely helps them establish a solid exercise habit, but as their need for external motivation decreases, they may feel less inclined to continue using these services.
@@ 3 Discussion
This study offers a detailed examination of the factors that influence the formation of consistent exercise habits among gym members. By applying advanced methodologies such as causal inference analysis and cluster-based segmentation, our findings provide nuanced insights that can inform the design of more effective interventions to promote regular physical activity.
One of the key distinctions of our study lies in the development and application of a survival metric to define and measure habit formation. This metric allowed us to capture the persistence of gym attendance over time more precisely than traditional methods. Previous studies that primarily relied on gym equipment data [25], here we offer a survival metric that identifies critical milestones in the early weeks of gym membership specifically, the 6-week and 17-week marks where the likelihood of maintaining regular attendance significantly declines using comprehensive visitation data. This approach provided a structured framework for understanding when and how gym habits are most vulnerable, thus offering clear targets for intervention.
Furthermore, the causal inference analysis in our study marked a significant advancement in understanding the causal impact of specific gym-related interventions on habit formation. Traditional observational studies often encountered challenges with confounding factors, making it difficult to accurately determine the true effects of various interventions [20, 17]. In contrast, our use of propensity score matching allowed us to more effectively isolate the effects of interventions such as group lessons, personal training sessions, visiting different clubs, and invitation credit usage. The analysis revealed that while interventions like personal training consistently demonstrated strong and sustained positive effects on habit formation, other interventions such as group lessons showed diminishing returns over time. However, it is important to highlight that interventions such as visiting different clubs and the use of invitation credits exhibited a more stable impact, with little decline over time. This suggests that while some social and motivational interventions may have temporally limited effectiveness, personalized and varied experiences like those offered through personal training and the opportunity to explore different gym environments maintain their influence over a longer duration. This finding shows the importance of offering a diverse range of interventions to cater to different aspects of habit formation, a detail that has been underexplored in previous studies, particularly those focusing on smaller social exercise settings [18].
@@
Building on the insights gained from this study, several avenues for future research and practical implementation can be explored. Firstly, this research aligns with several United Nations Sustainable Development Goals (SDGs) [27], particularly SDG 3 (Good Health and Well-being) and SDG 11 (Sustainable Cities and Communities). Promoting consistent exercise habits contributes to improved public health and well-being, supporting the development of healthier and more resilient communities. Encouraging sustained exercise routines, fitness centers, and public health initiatives can help reduce the incidence of these diseases, contributing to SDG 3’s target of reducing premature mortality from NCDs by one-third by 2030 [27].
Secondly, integrating wearable technology and fitness-tracking apps can provide real-time feedback and support a sense of accountability, which are critical components in sustaining exercise habits. Studies have shown that real-time feedback from wearable devices can significantly boost motivation and adherence to physical activity programs by reinforcing positive behaviors and helping users set realistic goals [28]. By providing continuous monitoring and personalized insights, these digital tools can help individuals stay on track with their fitness routines, adjusting their goals and strategies as needed to maintain engagement over time. This approach could be particularly effective when integrated with existing gym-based interventions, offering a hybrid model that utilizes both the precision of digital health tools and the motivational support of personal trainers.
@@
Finally, developing comprehensive intervention models that integrate psychological, social, and environmental factors can offer a more holistic approach to promoting regular exercise. Collaborating with behavioral scientists and public health experts can enhance the effectiveness of these models, ensuring that they address the multi-dimensional nature of habit formation. Additionally, machine learning can identify context variables associated with habit formation, which can inform targeted interventions to increase gym attendance, thereby optimizing the personalization and efficacy of these interventions [29].

### v2

@@ 3.1 Clustering Based on Visits
Figure 2: Behavioral clusters and preferences by demographics. NMF reveals 5 distinct exercise groups. These groups differ mainly by their preference to exercise during the day (a). Some of these clusters are more preferred by some gender and age groups (b). Since the cluster assignments were determined by user activities in the first 6 weeks, we measured the confidence of the NMF model for each cluster (c,d). Finally, we check the consistency of assignment when data for the first 6 and 17 weeks are considered (e).
Clustering analysis revealed five distinct clusters across all clubs, as shown in Figure 2(a). We identified five clusters corresponding to distinct daily time slots: morning (08:00–09:00), noon (12:00–13:00), afternoon (15:00–17:00), evening (18:00–19:00), and night (20:00–21:00). The morning and night clusters also exhibited distinct weekend patterns, which reflects the reduced weekend opening hours (08:00–20:00) compared to weekdays (06:00–23:00).
These behavioral groups were labeled according to their peak visit hours. The densities of the visits presented in the heatmaps reflect the most preferred times to visit the gym for each cluster. For this analysis, clustering was performed using only the first six weeks of visit data for each customer; however, we later demonstrate that the assigned clusters remain largely consistent when compared to visits through the $\text{17}^{\text{th}}$ week.
We then defined cluster-level deviations as in Section 2.3. With this definition, we wanted to see how the clusters deviated with respect to different variables we could group the customers, such as their gender and age groups. The deviation measures how much more or less likely it is to observe particular type of users within that cluster. For example, a deviation of 0.01 indicates that women are 1% more likely to appear in the Morning cluster compared to the overall population, as presented in the first subfigure in Fig2(b). Conversely, individuals aged 14–20 are 53% less likely to be found in the Morning cluster, as depicted in the same subfigure.
@@
Younger people (ages 14-27) tend to make fewer visits in the morning and noon, as indicated by their negative deviation in the two upper subfigures in Fig2(b). Conversely, older individuals (ages 49+) tend to visit more during the earlier parts of the day, as their morning deviations are positive while the rest are negative. This may be attributed to a preference for visiting before work to free up time afterward. The larger magnitude of deviations partly reflects the smaller sample size of older members. Women are more prominent in the noon and evening clusters, while men, as they make up the majority of people, do not deviate too much, but visit more often in the afternoon and at night. Most starkly, the people aged 14-20 make most of their visits in the afternoon. This likely reflects their school or university schedules, which permit afternoon gym attendance.
@@
Figures 2(c) and (d) show the cumulative distribution function for the probability of belonging to a cluster for each different cluster, signified by their unique color. These figures show that the NMF decomposition assigns morning and night clusters with higher confidence, whereas assignments for the intermediate clusters were less confident. This is also evident by the heatmaps produced in Fig2(a), as there is substantial overlap for the rest of the clusters. We hypothesize that individuals in the morning and night clusters may have more fixed schedules, which prevents them from visiting at other times. The rest of the clusters have more flexible schedules, allowing them to change their gym hours more frequently, thus making them harder to assign to a single cluster.
Figure 2(e) presents the cluster transition matrix, where the diagonal values represent the percentage of the population that remained in the same cluster between weeks 6 and 17. Although the previous two figures showed that morning and night clusters were more confidently assigned, this figure demonstrates that cluster assignments are robust over time. The diagonal values, which account for more than 84% of the population, indicate that members generally do not transition between different behavioral clusters. Additionally, most of the transitions are between clusters with adjacent time slots, which is to be expected. Thus, even though assignment is less confident for intermediate clusters, the observed stability suggests that underlying temporal preferences are persistent traits rather than random fluctuations. In sum, these clusters capture persistent, demographically patterned exercise routines. Their stability over time provides a robust foundation for subsequent causal inference and intervention analyses.
3.2 Identifying Critical Milestones for Habit Formation
@@
Our analysis of the survival metric revealed clear patterns in gym attendance behavior and allowed us to identify essential milestones that indicate the progression of habit formation. As shown in Figure 3(a), the cumulative distribution of members’ survival series shows a rapid increase in gym attendance in the first weeks. This rapid increase may reflect high initial motivation. To structure our interpretation, we aimed to define key milestones in this decline. Specifically, we identified the weeks at which 50% and 80% of users had ceased to maintain their attendance streaks. These thresholds were not selected based on any external theoretical rationale but served as pragmatic markers to summarize habit formation and dropout trends. Using this approach, we observed that approximately 50% of members did not maintain their attendance streak at the 6-week point. This 6-week milestone emerged as a critical point in the habit formation process where members either solidify or begin to abandon their gym habit. When we extended our analysis to users who survived even longer, we observed a steady decline in the number of members maintaining their attendance streak. At the 17-week point, 80% of the member population failed to maintain their streak, indicating an increasing difficulty in maintaining long-term attendance. Furthermore, our analysis showed that to pass this critical 6-week milestone, a member typically needed to make at least nine visits, establishing an average frequency of roughly two visits per week as a key behavior for sustained attendance (Figure 4(d)).
@@ 3.3 Impact of Interventions on Habit Formation
Figure 5(a) shows the overall effects of various gym-related interventions on habit formation.
@@
Personal trainer sessions exhibit a strong positive impact on maintaining regular exercise habits. The coefficients for personal training are consistently higher than those for other interventions, indicating the significant role of personalized guidance and accountability in habit formation. Personal trainers offer customized workout plans and motivational support, which appear to be significant in helping members establish and maintain their exercise routines.
@@
The results of the cluster analysis also indicate that specific interventions have diverse impacts on distinct subgroups of gym members. The impact of these interventions across different clusters, specifically highlighting the effects observed at the $\text{6}^{\text{th}}$ and $\text{17}^{\text{th}}$ weeks shown in Fig5(b).
@@
Noon cluster members exhibit a significantly higher positive response to invitation credit usage and group lessons compared to other clusters. This group tends to be older adults, with a higher-than-average proportion of females. Since the noon cluster hours coincide with working hours in Türkiye, it is possible that noon cluster members have more flexible schedules. The flexible schedules of the noon cluster may allow them to view the gym as a social hub, where they can combine exercise with socializing, which is particularly appealing to older adult members. Additionally, these members show a sustained positive effect from taking personal trainer sessions, particularly by week 17. This indicates that personalized guidance and accountability provided by personal trainers continue to support consistent exercise habits over time, making it an important intervention for this group.
@@
Evening cluster members have the highest proportion of females, which influences their response to interventions. Unlike other clusters, this group shows a lower positive response to group lessons. Although the female presence is significantly higher, similar to the noon cluster, the cluster behavior differs significantly. The overall effect of group lessons on the evening cluster is negligible, with the impact being lower than the entire population (0.18), suggesting that group workouts may not be as appealing to them. However, this cluster shows the highest and most sustained positive effect from personal trainer sessions by the $\text{17}^{\text{th}}$ week, making it the cluster that best maintains its benefits from personalized training. This indicates that for evening exercisers, particularly in a female-represented cluster, personalized attention and tailored workouts play a crucial role in fostering long-term exercise commitment, especially during the evening hours after a day’s work.
Night cluster benefits from a mix of interventions, including group lessons and personal training sessions. The late-night hours are typically suitable for those with unconventional schedules, such as students or shift workers. The fact that this group is predominantly younger, with older individuals being strongly underrepresented also supports this. Working with a personal trainer has the most significant effect on habit formation within this cluster during the first 6 weeks. However, this effect diminishes considerably by the $\text{17}^{\text{th}}$ week. This decline may suggest that younger members, who form the majority of this cluster, initially benefit from the structure and guidance provided by a personal trainer but may choose to transition to more independent workouts as they gain confidence and familiarity with their routines. The initial high impact of personal training likely helps them establish a solid exercise habit, but as their need for external motivation decreases, they may feel less inclined to continue using these services.
@@ 4 Discussion
This study offers a detailed examination of the factors that influence the formation of consistent exercise habits among gym members. By applying methodologies such as causal inference analysis and cluster-based segmentation, our findings provide nuanced insights that can inform the design of more effective interventions to promote regular physical activity.
One of the key distinctions of our study lies in the development and application of a survival metric to define and measure habit formation. This metric allowed us to capture the persistence of gym attendance over time more precisely than traditional methods. Previous studies primarily relied on gym equipment data (Harris and Kessler, 2019). In contrast, here we offer a survival metric that identifies critical milestones in the early weeks of gym membership specifically, the 6-week and 17-week marks where the likelihood of maintaining regular attendance significantly declines using comprehensive visitation data. This approach provided a structured framework for understanding when and how gym habits are most vulnerable, thus offering clear targets for intervention.
Furthermore, the causal inference analysis in our study marked a significant advancement in understanding the causal impact of specific gym-related interventions on habit formation. Traditional observational studies often encountered challenges with confounding factors, making it difficult to accurately determine the true effects of various interventions (Fischer and Bryant, 2008; Burke et al., 2005). In contrast, our use of propensity score matching (PSM) allowed us to more effectively isolate the effects of interventions such as group lessons, personal training sessions, visiting different clubs, and invitation credit usage. The analysis revealed that while interventions like personal training consistently demonstrated strong and sustained positive effects on habit formation, other interventions such as group lessons showed diminishing returns over time. However, it is important to highlight that interventions such as visiting different clubs and the use of invitation credits exhibited a more stable impact, with little decline over time. This suggests that while some social and motivational interventions may have temporally limited effectiveness, personalized and varied experiences like those offered through personal training and the opportunity to explore different gym environments maintain their influence over a longer duration. This finding shows the importance of offering a diverse range of interventions to cater to different aspects of habit formation, a detail that has been underexplored in previous studies, particularly those focusing on smaller social exercise settings (Wayment and McDonald, 2017).
@@
Building on the insights gained from this study, several avenues for future research and practical implementation can be explored. Firstly, this research aligns with several United Nations Sustainable Development Goals (SDGs) (Nations, 2015), particularly SDG 3 (Good Health and Well-being) and SDG 11 (Sustainable Cities and Communities). Promoting consistent exercise habits contributes to improved public health and well-being, supporting the development of healthier and more resilient communities. Encouraging sustained exercise routines, fitness centers, and public health initiatives can help reduce the incidence of these diseases, contributing to SDG 3’s target of reducing premature mortality from NCDs by one-third by 2030 (Nations, 2015).
Secondly, integrating wearable technology and fitness-tracking apps can provide real-time feedback and support a sense of accountability, which are critical components in sustaining exercise habits. Studies have shown that real-time feedback from wearable devices can significantly boost motivation and adherence to physical activity programs by reinforcing positive behaviors and helping users set realistic goals (Patel et al., 2015). By providing continuous monitoring and personalized insights, these digital tools can help individuals stay on track with their fitness routines, adjusting their goals and strategies as needed to maintain engagement over time. This approach could be particularly effective when integrated with existing gym-based interventions, offering a hybrid model that utilizes both the precision of digital health tools and the motivational support of personal trainers.
@@
Finally, developing comprehensive intervention models that integrate psychological, social, and environmental factors can offer a more holistic approach to promoting regular exercise. Collaborating with behavioral scientists and public health experts can enhance the effectiveness of these models, ensuring that they address the multi-dimensional nature of habit formation. Additionally, machine learning can identify context variables associated with habit formation, which can inform targeted interventions to increase gym attendance, thereby optimizing the personalization and efficacy of these interventions (Buyalskaya et al., 2023).